### v1

# Embodied Crowd Counting

Abstract
Occlusion is one of the fundamental challenges in crowd counting. In the community, various data-driven approaches have been developed to address this issue, yet their effectiveness is limited. This is mainly because most existing crowd counting datasets on which the methods are trained are based on passive cameras, restricting their ability to fully sense the environment. Recently, embodied navigation methods have shown significant potential in precise object detection in interactive scenes. These methods incorporate active camera settings, holding promise in addressing the fundamental issues in crowd counting. However, most existing methods are designed for indoor navigation, showing unknown performance in analyzing complex object distribution in large scale scenes, such as crowds. Besides, most existing embodied navigation datasets are indoor scenes with limited scale and object quantity, preventing them from being introduced into dense crowd analysis. Based on this, a novel task, Embodied Crowd Counting (ECC), is proposed. We first build up an interactive simulator, Embodied Crowd Counting Dataset (ECCD), which enables large scale scenes and large object quantity. A prior probability distribution that approximates realistic crowd distribution is introduced to generate crowds. Then, a zero-shot navigation method (ZECC) is proposed. This method contains a MLLM driven coarse-to-fine navigation mechanism, enabling active Z-axis exploration, and a normal-line-based crowd distribution analysis method for fine counting. Experimental results against baselines show that the proposed method achieves the best trade-off between counting accuracy and navigation cost.
Figure 1: Comparisons between our dataset&model and others. (a) Comparison of environment areas between existing robot navigation datasets and ours. Our dataset features large-scale outdoor crowd scenes. (b) Comparison of lenses interactivity between existing crowd counting datasets and ours. Our dataset includes a 3D structural environment for each scenario, enabling interaction with agents. (c) Comparison of methods between traditional crowd counting models and ours. We propose using agent navigation to achieve precise detection in previously unexplored faraway areas.
[…]
1 Introduction
Crowd counting is critical for public safety and urban planning [21]. One main challenge in this field is occlusion. It can be categorized into two aspects: overlap and invisibility. Overlap refers to the high density of people stacked together, making it difficult to distinguish each individual from some viewpoints, while invisibility indicates that the current camera position is obstructed, such as being blocked by buildings, or far from the crowds so that the target cannot be detected clearly. To summarize, these situations are caused by a poor observation point. Existing methods try to solve these challenges from several perspectives, such as using multi-scale feature extraction [20], body part detection [26] or multi-camera fusion [49, 50]. Datasets are also expanded for enhancing method performance [51, 19, 36, 43]. In general, these methods or datasets either try to enhance learned model representation, or expand the camera sensing range by introducing multi-view settings. Yet, their passive camera settings do not fully solve the occlusion challenge in crowd counting, especially in cases where crowds exceed the sensing range, or no cameras are set to detect completely obscured crowds. These issues restrict passive-camera-based methods in practice.
Recent development in Embodied AI brings perspectives to solve the occlusion challenge in crowd counting. It has been demonstrated to possess significant potential in enhancing scene exploration and object detection. Aspects such as Vision-Language-Navigation (VLN) have been directed towards equipping mobile robots with human-like perception abilities, resulting in remarkable performance in exploring scenes and detecting objects in open environments [9, 4, 52, 7]. The active camera settings in VLN is promising to solve the fundamental challenges in crowd counting, since it can optimize observation point to mitigate overlap and invisibility caused by poor camera position setting. Yet most VLN benchmarks [44, 40, 3, 1] are indoor environments with limited exploration space (e.g., no Z-axis action options) and relatively small object quantity. And the performance of such methods remains unknown for detecting crowds with a large quantity and complex distribution in large-scale scenes. This results in a significant gap between VLN and crowd counting, as in practice, crowds often appear in large spaces with varying distribution.
To address the issues mentioned above, we first define a novel task, Embodied Crowd Counting (ECC). This task integrates crowd counting into the Zero-shot Object-Goal Navigation task (ZSON) in VLN [28]. It is designed for large-scale scenes with a large object quantity. A testing environment, Embodied Crowd Counting Dataset (ECCD), is then developed for this task. Compared to existing crowd counting and VLN datasets, this test environment has the following features: 1) It supports 60 distinct and diverse large-scale outdoor scenes. Each scene covers an area up to 40,000 $m^{2}$, with target quantity up to 15,000. 2) It uses a prior distribution, Poisson Point Process [11], for crowd quantity distribution modeling, which approximates realistic crowd situations.
Based on this platform, we develop a method, Zero-shot Embodied Crowd Counting (ZECC), to conduct crowd counting in ECCD. Compared to existing crowd counting methods and VLN methods, ZECC has the following features: 1) It uses active camera setting to solve fundamental challenges in crowd counting, transferring the detection task to an exploration task; 2) It consists of a coarse-to-fine navigation mechanism which uses common sense of Multi-Modal-Large-Language-Model (MLLM) to analyze the surrounding environment and plan Z-axis exploration, leveraging advantages of 6 degrees of freedom (6-DoF) to avoid low-altitude obstacles and gain border sight, thus achieving efficient exploration; 3) It consists of a crowd distribution analysis mechanism which transfers the complex crowd distribution to simple surfaces. The normal lines of the surfaces are used to generate fine observation points, which are used to mitigate the influence of overlap and invisibility, enabling accurate crowd counting.
The contributions of this work can be summarized as follows:
  • We present an innovative task called ECC, specifically designed to address the challenges of occlusion and multi-scale complexities that are prevalent in conventional crowd counting methods.
  • A new dataset called ECCD has been collected to redefine the landscape of crowd analysis. Unlike traditional crowd counting and VLN datasets, it features large-scale outdoor crowd scenes with interactive capabilities.
  • ZECC is proposed to use MLLM for Z-axis exploration, reducing costs while ensuring detection performance. By utilizing normal lines to calculate navigation points, this approach eliminates overlap and enhances visibility in crowds.
  • Experiments show that the proposed method outperforms the baseline methods in various testing environments.
2 Related Work
2.1 Crowd Counting
Crowd-counting algorithms have greatly benefited from large-scale, high-quality datasets like UCF-CC50 [18] and UCF-QNRF [19]. These foundational datasets have facilitated the creation of subsequent collections focused on dense crowd imagery, such as ShanghaiTech [51], JHU-CROWD++ [36], NWPU-Crowd [43]. However, the images in these datasets are generated from fixed cameras. In contrast, ECCD provides interactive capabilities while maintains diverse crowd distribution.
Methods like [51, 20, 24, 24, 5] leverage crowd distribution prior in an image, or using attention maps to learn dependency. Recent multi-modal approaches [10, 47, 29, 42] leverage vision-language models to transfer image-text knowledge to dense crowd prediction. While these models improve long range and overlapping small target detection ability, their performance are restricted in practical if overlap reaches extreme level. Recent efforts expand spatial coverage. These works include multi-view systems [32, 17] that use multi camera to capture images from larger scale scenes. Others like [14, 15] use recorded video to conduct crowd analysis. While these are important advances on solving the fundamental challenges in crowd counting against the single-fix-image-based methods, their camera setting is still predefined by the dataset and cannot be adjusted when inferring. This cannot ensure fully exploration in a large environment. ZECC allows active exploration, which is fundamentally different from existing methods.
2.2 Embodied Navigation
Many traditional robot navigation methods were developed using conventional navigation datasets like KITTI [12] and SUN RGB-D [38]. Beyond these foundational datasets, [44, 33] provide 3D indoor environments for navigation and interaction tasks. These datasets are limited with their small scale size and small object quantity. Recent datasets have been created towards outdoor navigation like AerialVLN [25], and CityNav [23]. However, they are designed for VLN tasks without consideration of object quantity. Compared to these datasets, ECCD simultaneously supports large scale outdoor scenes and large object quantity with diverse distribution.
Efficient exploration using mobile robots remains a crucial challenge in vision and robotics. [13, 16, 30, 38] have developed human-like cognitive maps, enabling autonomous path learning in unknown environments. Other approaches like [2, 6, 31] use reinforcement learning to develop exploration policies. Recently, [48] applied video-based visual-language models to plan sequential actions in VLN. And [27, 53] presented zero-shot models that utilizes natural language instructions to guide agents through environments without prior environment-specific training. However, the methods mentioned above are for indoors navigation without Z-axis moving ability. This restricts their ability in large scale outdoor scenes. Recently, methods like [25, 45] introduce 6-DoF in outdoor VLN. Yet they are designed for instruction following task, in which the movement is restricted by human language. Besides, they lack ability to analyze crowds with large quantity and complex distribution. In contrast, ZECC is the first self-motivated agent in 6-DoF and can handle crowds simultaneously.
Figure 2: The proposed framework for ZECC consists of several key steps. First, we utilize an exploration strategy based on FBE to create a global crowd distribution by generating a target point cloud. Next, we apply a Gaussian Mixture Model (GMM) clustering to identify the representative centers of the global target point cloud and determine the final navigation points based on potential navigation directions. Finally, we integrate detection results obtained from various detectors with depth observations, mapping these to the global point cloud. The number of filtered targets is then counted and recorded as the final result for the crowd.
3 Method
3.1 Problem Definition
To ensure the interactive with environment for accurately crowd counting in vast outdoor environments, we propose an innovative task called Embodied Crowd Counting (ECC). In ECC, an agent is first deployed in an unknown environment. At time step $t$, the input of the agent is RGB-D observations ${O_{t}}$ along with its pose ${p_{t}}$. Based on this information, the agent predicts navigation point $p\subset\mathbb{R}^{3}$ to a drone at time step $t$ and the drone moves to $p$. During exploration, the agent is allowed to record observations that help with crowd counting. When the agent decides to stop, it outputs an integer that represents its crowd counting result. The counting error and travel distance are calculated to assess the agent’s performance. ECC can be considered analogous to the ZSON task, since they both require an agent to detect targets in an unexplored environment without additional assistance. However, there are several differences. The task setting differences are: 1) The object category $G$ in ZSON is limited to crowd in ECC. 2) The step limitation in ZSON is removed in ECC. 3) The evaluation criterion in ECC is changed to counting error and travel distance. These differences do not prevent the availability of most recent ZSON methods in ECC by conducting minor modifications. However, agents should face essential differences in ECC: 1) Z-axis is available, which is not considered by most recent ZSON methods. 2) Complex crowd distributions exist, including heavy occlusions, while ZSON methods only consider objects with small quantity. Current ZSON methods show unknown performance under such differences, and new methods should be considered in ECC. Note that ECC is different from the instruction following tasks in VLN such as [25, 45], since these tasks require natural language assistance to drive the agent.
3.2 Embodied Crowd Counting Dataset (ECCD)
We propose a new dataset called ECCD to support algorithm design and evaluation for the community. ECCD is developed using Unreal Engine 4. This platform enables programmable environment construction, allowing scene richness and scalability. The characteristics of ECCD are as follows:
Diversity. Since different environments have different layouts, crowd distributions, and quantities reflect different challenges, ECCD is designed to contain diverse scenes. This dataset contains 60 distinct environments. Also, ECCD has an area up to 40,000 $m^{2}$, reaches a height up to 50 $m$, and allows for the simultaneous presence of over 15,000 targets within a single scene. This is significantly different from existing robot navigation datasets, such as [33] which offers environments with an average navigable space of 1,000 $m^{2}$, and crowd counting datasets, such as [51] which contains 1,198 annotated images captured with static cameras and a total of 330,165 individuals.
Realism. The ECCD is designed to simulate large-scale outdoor crowd scenarios in reality, in order to ensure practical effectiveness of systems built on ECCD, as shown in Figure 3. Environments like a city, a stadium, and a parking lot are simulated. Additionally, the environments support a complex structure of buildings in real life, such as multiple floors and bleachers. These features allow ECCD to reflect challenges in reality.
Figure 3: ECCD is designed to realistically mimic building and crowd distribution. On the left are samples from ECCD, and on the right are the real scenes.
Crowd generation mechanism. To model real challenges in crowd counting, ECCD uses Poisson Point Process for crowd quantity distribution modeling [11]. In ECCD, Blocks are placed in the simulator to present potential areas where crowds may exist. Then, for each block $\mathcal{U}\subset\mathbb{R}^{2}$, the process is defined as:
$\mathbb{P}\left(N(\mathcal{U})=k\right)=\frac{(\lambda\cdot|\mathcal{U}|)^{k}}{k!}e^{-\lambda\cdot|\mathcal{U}|},\quad k\in\mathbb{N},$ (1)
where $N(\mathcal{U})$ denotes the number of individuals in block $\mathcal{U}$, $\lambda$ represents the crowd density set by human experts according to the environment semantics, and $|\mathcal{U}|$ is the area of the block. This ensures ECCD generates crowds that approximate real situations. By comparison, existing simulators based on UE4, such as AirVLN and OpenUAV, do not consider object quantity and distribution.
3.3 Zero-shot Embodied Crowd Counting (ZECC)
3.3.1 Overview
Previous embodied navigation agents are designed for indoor environments [27], or relying on language assistance [25, 45], making them restricted in large-scale outdoor scenes. Under such context, a zero-shot agent, ZECC, which can actively control altitude and conduct crowd analysis, is proposed. As illustrated in Figure 2, the method consists of 2 components, Active Top-down Exploration (ATE) and Normal-line based Navigation (NLBN). ATE is designed for adjusting agent altitude for efficient coarse crowd distribution estimation, and NLBN estimates normal lines on top of crowd for accurate crowd observation, alleviating occlusion.
3.3.2 Active Top-down Exploration (ATE)
Crowds are often concentrated in specific areas, such as roads and squares, making it unnecessary to explore every region of the environment. To address this, ATE is proposed to estimate the global distribution of crowds from a high altitude. This approach aims to improve efficiency by focusing exploration efforts on the most relevant areas and avoiding unnecessary locations. In particular. high-altitude exploration (HAE) brings a broader field of view for decision making, as well as relatively less exploration cost since obstacles are often sparse in high altitude. While low-altitude exploration (LAE) provides close range observation for clear target detection. It is essential for the agent to plan HAE and LAE to achieve efficiency and accuracy simultaneously. Therefore, ATE leverages the Z-axis mobility of outdoor agents and the common sense of multi-modal large language model (MLLM) to switch between HAE and LAE by local environment layout reasoning. Then, the crowd distribution is estimated by crowd counting model to predict density maps in the environment.
Specifically, the agent collects observations ${O_{t}}=\left\{{o_{t}^{1},\ldots,o_{t}^{c}}\right\}$ and pose $p_{t}$ at time step $t$, where $c$ indicates the camera number of the agent. Then, a MLLM is prompted by using the observations and text prompts to conduct environment layout reasoning. The MLLM is asked to predict whether current location is valuable for LAE, by refering current crowd appearence and obstacle layout. This process is formulate as:
$s_{t}=\text{MLLM}\left({{O_{t}};I}\right),$ (2)
where $\text{MLLM}\left(\cdot\right)$ is the inference process and $I$ is the prompt. $s_{t}\in\left[{0,1}\right]$. If $s_{t}>0.5$, the agent will adjust its altitude for LAE. After switching altitude strategy, the agent will conduct regular exploration. For HAE, one of the frontiers between explored and unknown areas is selected as navigation point. For LAE, the agent will keep exploring frontiers until areas within its field of view during HAE are fully explored. During LAE, once the agent reaches a frontier $f$, it gathers observations ${O_{f}}=\left\{{o_{f}^{1},\ldots,o_{f}^{c}}\right\}$, and a crowd counting model predicts crowd density maps on the observations. Then the density maps are projected onto the global point cloud to form a global crowd distribution. This is formulated as:
$d_{f}=\text{P}\left({\text{G}\left({{O_{f}}}\right),{p_{f}}}\right),$ (3)
where $\text{P}\left(\cdot\right)$ is the projection operation, $\text{G}\left(\cdot\right)$ is a crowd counting model, and $p_{f}$ is the agent pose at $f$.
3.3.3 Normal-line based Navigation (NLBN)
NLBN is utilized for the precise analysis of potential overlapping structures within crowds. The intuition behind this approach is that while a random viewpoint may often obscure visibility in a dense crowd, there are several vantage points that can still provide a clear view for identifying individuals. These observation points are expected to be above and maintain a certain degree to the center of the crowd, since a top-down view can distinguish overlapping, and perspectives from certain angles can ensure that the targets are in the field of view (FOV) of the agent. To find such observation points, NLBN transfers the crowd detection task into a surface detection task. By fitting surfaces on the crowd distribution, normal lines of the surfaces can be obtained, and view points which fulfill the constraints mentioned above can be accurately calculated by using normal lines.
After achieving global crowd distribution in ATE, NLBN first clusters the crowd point cloud into subregions using a clustering method. This step helps divide the large crowd point cloud into manageable parts, as navigating to each navigation point can be resource-intensive. Next, surfaces are fitted to these subregions based on the point cloud data to obtain normal lines. This process transforms the challenging task of crowd analysis into a more manageable surface analysis. Finally, optimized navigation points are generated based on the normal lines, ensuring they maintain a certain distance and angle from the crowd surfaces for accurate crowd detection. A viewpoint-based method is employed to confirm that the selected navigation points do not overlap with any obstacles.
In particular, GMM [35] is used to divide the global crowd distribution into manageable subregions. It ensures each sample point in a distribution can be assigned with a cluster. Thus, it can divide any crowd distribution into patches. A parameter $\epsilon$ is used to determine the size of each GMM cluster.
After the clusters and centers are generated, surfaces are fitted for each patch. Then, for the $N$-th cluster, the normal is obtained and represented by ${\bf{d}}_{N}^{{\rm{cluster}}}$. Candidate view directions $\{\mathbf{d}_{N1}^{\text{view}},...,\mathbf{d}_{Nm}^{\text{view}}\}_{N}$ are sampled under angular constraints:
$\frac{\mathbf{d}_{N}^{\text{cluster}}\cdot\mathbf{d}_{Nm}^{\text{view}}}{\|\mathbf{d}_{N}^{\text{cluster}}\|\|\mathbf{d}_{Nm}^{\text{view}}\|}=\zeta,$ (4)
where $\zeta$ is a hyper parameter. These candidate view directions can bring the agent to a optimized observation point by selecting a position along the vector. However, there are two issues: 1) The position may be located on obstacles; 2) The crowd cluster may be out of the agent’s FOV. To solve the issues and identify the final navigation points from the potential ones, a simple yet effective technique is proposed. In the process of ATE, each cluster point identified by the counting model processes at least one viewpoint of the agent. The vector connecting the viewpoints and target points firmly indicates navigable space, since the propagation of light naturally points to unobstructed area. Besides, the vector can ensure that the crowd is in the FOV as well, since the target image is identified during ATE. Based on this prior, the ATE viewpoints are used to generate the final navigation points. The ATE view vector is calculated as:
$\mathbf{\bf{d}}_{N}^{{\rm{ATE}}}={\bf{x}}_{N}^{{\rm{ATE}}}-{\bf{x}}_{N}^{{\rm{cluster}}},$ (5)
where ${\bf{x}}_{N}^{{\rm{ATE}}}$ is the navigation point from which the agent finds the cluster center ${\bf{x}}_{N}^{{\rm{cluster}}}$ during ATE. Then, the potential navigation directions that are at the minimum angles to the ATE view vectors are selected as the final navigation directions:
$\mathbf{d}_{N}^{\text{view}}=\mathop{\arg\min}\limits_{m}\frac{\mathbf{d}_{N}^{\text{ATE}}\cdot\mathbf{d}_{Nm}^{\text{view}}}{\|\mathbf{d}_{N}^{\text{ATE}}\|\|\mathbf{d}_{Nm}^{\text{view}}\|},$ (6)
and the final navigation point $\mathbf{x}_{N}^{\text{view}}$ is calculated by
$\mathbf{x}_{N}^{\text{view}}=\mathbf{x}_{N}^{\text{cluster}}+\eta\cdot\mathbf{d}_{N}^{\text{view}},$ (7)
where $\eta$ is a hyper parameter representing the distance between the agent and the crowd cluster. NLBN ensures close-range, precise target observation and safe navigation, even in complex and unfamiliar environments. The potential navigation vectors, normal lines, and ATE view vectors are illustrated in Figure 4.
Figure 4: Illustration of the potential navigation vectors, normal lines, and FBE view vectors. Zoom in for better visualization.
3.3.4 Fine Detection and Counting (FDC)
Using the navigation points generated by NBLN, the agent travels from one point to another through path planning algorithms. Upon reaching each navigation point, close-range and high-resolution RGB observations can be conducted. These observations are then utilized as inputs for detection models to perform precise target detection. The detection results are projected onto the global point cloud using the depth sensor and the agent pose. To prevent repetitive detections, only one target is retained for each region within a specific scale. Ultimately, the result is calculated by counting the quantity of filtered targets.
4 Experiments
4.1 Settings
Metrics. Mean Absolute Percentage Error (MAPE) is used as the metric to evaluate the counting performance, which is defined as:
[…]
where $M$ is the quantity of testing environments, $\hat{y}_{i}$ and $y_{i}$ are the estimated count and the ground truth count in an environment, respectively.
The sum of Euclidean distance between adjacent navigation points along the agent traveling path is used to evaluate the travel distance TD, which is defined as:
${\rm{TD}}=\sum\limits_{i=1}^{n-1}{\left\|{{{\bf{x}}_{i+1}}-{{\bf{x}}_{i}}}\right\|},$ (9)
where $n$ is the quantity of navigation points in one episode, ${{{\bf{x}}_{i}}}$ and ${{{\bf{x}}_{i+1}}}$ are the coordinates of the two adjacent navigation points, respectively.
Implementation Details. MLLM used in ATE is GPT-4V [46]. The path planning method is A* algorithm [37]. Altitude is $80m$ for HAE and $10m$ for LAE. The crowd density estimator in ATE is Generalized Loss (GL) [41], and the detection model in FDC is Grounding DINO (GD) [34]. For hyper parameters, navigation vector deg $\zeta$ is 15 °, density threshold $\kappa$ is 0.7, navigation point range $\eta$ is 8 m, and cluster size $\epsilon$ is 40.
Baselines. We compare ZECC with existing exploration method Frontier-based exploration (FBE) [39], as well as SOTA ZSON methods, 1) CoW [8], 2) OpenFMNav [22]. These methods use advanced foundation models to conduct decision making during navigation. To apply the baselines in ECC, the exploration step limitation for ZSON methods is removed, and they are equipped with GD or GL to conduct crowd counting using the projection method in FDC.
4.2 Results
Overall Performance. We report overall performance in Table 1. ZECC offers the optimal balance between counting performance and cost. In contrast, FBE delivers the best TD but lacks the ability to perform a target detection function, which limits its capacity for fine observation. This incomplete observation negatively impacts its overall performance. ZSON methods, such as Cow and OpenFMNav come with perception modules; however, they do not select a 3D fine observation point as NLBN, which means they cannot fully observe each individual in the crowd. This limitation leads to individuals being obscured by overlapping crowds, especially when the density is high. In some instances, these methods only identify single individuals, which reduces their efficiency. In contrast, ZECC improves counting efficiency by planning Z-axis exploration using ATE and employing NLBN for detailed detection.
For agents utilizing counting models for crowd counting, the performance tends to be slightly lower than that of agents using detection models. This is due to the increased sparsity of crowd observations as the agents get closer to the crowd. Crowd counting models are mainly trained on images depicting dense crowds, which results in limited generalization capability for scenarios with sparser crowds. Additionally, current ZSON methods do not consider finding a close position for crowd observation, which further degrades performance.
Method MAPE (%) TD (m)
FBE [39] + GL [41] 55.92 2227.71
FBE [39] + GD [34] 53.85 2227.71
CoW [8] + GL [41] 51.19 3271.07
CoW [8] + GD [34] 46.70 3271.07
OpenFMNav [22] + GL [41] 59.71 4936.79
OpenFMNav [22] + GD [34] 46.84 4936.79
ZECC 18.91 3804.63
Table 1: Overall performance of the methods. ZECC achieves best trade-off between MAPE and TD.
Figure 5: Performance and cost of ZECC and the baselines under different crowd density levels. L1-L5 refers to increasing density level. The figure demonstrate that ZECC achieves a balance between performance and exploration cost.
Figure 6: The effect of four hyperparameters used in our model, including GMM cluster size, navigation vector degree, navigation point range and density map threshold.
We analyze various environments with different crowd density levels and visualize the average performance and cost of the methods using the GD detector in these scenarios. The results are presented in Figure 5. Among the methods evaluated, ZECC demonstrates the best average performance across different density levels. As crowd density increases, the MLLM in ATE becomes more effective at identifying high-density crowds, resulting in improved performance. In contrast, other methods struggle to manage high-density crowds, leading to limited performance.
In terms of cost, ZECC’s navigation points are influenced by crowd distribution and density. As the density level increases, the cost also rises. Although ZECC falls short of achieving the lowest cost in the last two density levels, its costs are still comparable to the baseline while ensuring effective counting performance. In contrast, other methods do not actively adjust navigation points based on crowd distribution, which results in their costs remaining relatively stable and limits their effectiveness in low-density scenes.
MAPE (%) TD (m)
ZECC w/o ATE 17.46 4633.67
ZECC w/o HAE 17.46 4633.67
ZECC w/o LAE 88.08 1738.84
ZECC 18.91 3804.63
Table 2: The results of ablation study for components in ATE. w/o ATE refers to using FBE + NLBN results for crowd counting. w/o HAE refers to fix the agent’s altitude to LAE. w/o LAE refers to fix the agent’s altitude to HAE. ZECC improves approximately 17% cost by reducing approximately 8% performance, achieving a better trade-off.
MAPE (%) Successful rate (%)
ZECC w/o NLBN 65.19 100.00
ZECC w/o NL 98.44 8.33
ZECC w/o VPS 92.55 100.00
ZECC w/o ATE-VPS 99.49 1.45
ZECC 18.91 100.00
Table 3: The results of ablation study for components in NLBN. w/o NLBN refers to using ATE results for crowd counting. w/o NL refers to not using normal line (NL) to calculate navigation points but use the cluster centers as navigation points. w/o VPS refers to not using view point selection (VPS) but select a point along the normal vector with $\eta$. w/o ATE-VPS refers to not using the ATE-view-vector-based view point selection (ATE-VPS), but randomly select a navigation vector from the potential navigation vectors. Successful rate indicates the ratio of the reachable navigation point reported by the A* path planning algorithm.
4.3 Ablation Study
ATE. We conducted ablation studies by removing specific components from the ATE. The results are presented in Table 2. The configuration without ATE (w/o ATE) achieves the best performance because it explores the environments greedily and utilizes NLBN for precise detection, resulting in adequate perception. However, this approach leads to a larger temporal delay (TD). In contrast, ZECC achieves a better balance by conducting both HAE and LAE simultaneously. w/o HAE uses the same setting as w/o ATE. w/o LAE fails for it maintains at a high altitude and does not conduct close range detection, making it hard to capture targets.
NLBN. We then conducted ablation studies on NLBN and the results are presented in Table 3. The findings indicate that omitting any component of NLBN results in a significant drop in either performance or success rates. Without NLBN, the absence of optimized viewpoints causes ZECC to revert to a ZSON method. When both NL and FBE-VPS are removed, most navigation points end up being located on obstacles, leading to a scarcity of reachable navigation points and incomplete exploration. Additionally, without VPS, the navigation points are positioned directly above the targets, causing the targets to fall out of the field of view, which results in corrupted observations.
4.4 Hyper parameter Study
We then examine the effects of parameters in the method, including GMM cluster size, navigation vector degree, navigation point range, and density map threshold. A gym-like scene featuring a densely packed crowd is utilized to test these parameters. The results are illustrated in Figure 6.
GMM cluster size. As the size of the GMM cluster increases, performance tends to deteriorate. When the GMM cluster size is set too low, the method generates sparse navigation points, which prevents the agent from effectively exploring crowded areas, resulting in poor counting performance. Although a smaller GMM cluster size can initially yield high performance, the results become worse when the cluster sizes reach 30 and 40, respectively. Specifically, the cost for a cluster size of 30 is 4065.02 meters, while for a size of 40, it rises to 5125.44 meters. Opting for an extreme GMM cluster size may significantly increase costs without providing substantial improvements in performance. Therefore, it is crucial to carefully determine the GMM cluster size to find a balance between cost and performance.
Navigation vector degree. Performance is optimal at 20°, while it declines at other angles. This degradation occurs because the agent has a limited field of view; it cannot effectively observe targets outside this range. This highlights the importance of our navigation point generation method.
Navigation point range. The performance is optimal at a distance of 8 meters, while performance declines at both shorter and longer ranges. A similar trend is observed with the navigation vector degree. At extremely close range, the agent’s field of view is restricted, and at long range, the agent is unable to gather detailed observations, which negatively impacts the subsequent target detection phase. Therefore, selecting the right viewpoint and generating appropriate navigation points are crucial for effective scene exploration algorithms.
Density map threshold. The results indicate that similar to GMM cluster size, performance declines as the threshold increases. A lower density map threshold results in a coarser estimation of the target zone, which in turn generates more navigation points. This allows for more detailed exploration and observation, leading to improved counting performance. Yet cost rises. For instance, when the density map threshold is 0.7, the travel distance is 4720.02 meters, and it increase to 7272.02 meters when the threshold is 0.4. This is a significant degradation in cost, yet improvement in performance is minor. This emphasizes the importance of parameter selection.
5 Conclusion
In this study, Embodied Crowd Counting (ECC) is proposed, a task which enables interactive crowd counting. A simulator, the Embodied Crowd Counting Dataset (ECCD) is developed to enable related research for ECC. This dataset includes 60 diverse virtual environments with crowd density modeled by a prior probability distribution, approximating the reality. A method, Zero-shot Embodied Counting (ZECC), is proposed to verify this task. This is an active agent which can explore unknown environments without additional assistance. Active Top-down Exploration (ATE) is proposed to utilize Z-axis moving ability for exploration planning. This module is equipped with MLLM to active high altitude exploration (HAE) or low altitude exploration (LAE), balancing crowd counting performance and exploration cost. Normal-line based Navigation (NLBN) is proposed to select optimized navigation point for crowd observation. This module generates a navigation point from the top-down view and maintains an angle, alleviating overlap of the crowd. Simultaneously, the estimate navigation points enable obstacle avoidance and ensure that the crowd is in FOV. Experiment results show that ZECC achieves a balance between performance and cost compared to recent navigation agents. Future research will focus on more efficient and scalable agents to enhance performance and adaptability.

### v2

@@
 These methods incorporate active camera settings, holding promise in addressing the fundamental issues in crowd counting.
-However, most existing methods are designed for indoor navigation, showing unknown performance in analyzing complex object distribution in large scale scenes, such as crowds.
+However, most existing methods are designed for indoor navigation, showing unknown performance in analyzing complex object distribution in large-scale scenes, such as crowds.
 Besides, most existing embodied navigation datasets are indoor scenes with limited scale and object quantity, preventing them from being introduced into dense crowd analysis.
-Based on this, a novel task, Embodied Crowd Counting (ECC), is proposed.
-We first build up an interactive simulator, Embodied Crowd Counting Dataset (ECCD), which enables large scale scenes and large object quantity.
-A prior probability distribution that approximates realistic crowd distribution is introduced to generate crowds.
-Then, a zero-shot navigation method (ZECC) is proposed.
-This method contains a MLLM driven coarse-to-fine navigation mechanism, enabling active Z-axis exploration, and a normal-line-based crowd distribution analysis method for fine counting.
-Experimental results against baselines show that the proposed method achieves the best trade-off between counting accuracy and navigation cost.
-Figure 1: Comparisons between our dataset&model and others.
-(a) Comparison of environment areas between existing robot navigation datasets and ours.
-Our dataset features large-scale outdoor crowd scenes.
-(b) Comparison of lenses interactivity between existing crowd counting datasets and ours.
-Our dataset includes a 3D structural environment for each scenario, enabling interaction with agents.
-(c) Comparison of methods between traditional crowd counting models and ours.
-We propose using agent navigation to achieve precise detection in previously unexplored faraway areas.
-[…]
+Based on this, a novel task, Embodied Crowd Counting (ECC), is proposed to count the number of persons in a large-scale scene actively.
+We then build up an interactive simulator, the Embodied Crowd Counting Dataset (ECCD), which enables large-scale scenes and large object quantities.
+A prior probability distribution approximating a realistic crowd distribution is introduced to generate crowds.
+Then, a zero-shot navigation method (ZECC) is proposed as a baseline.
+This method contains an MLLM-driven coarse-to-fine navigation mechanism, enabling active Z-axis exploration, and a normal-line-based crowd distribution analysis method for fine counting.
+Experimental results show that the proposed method achieves the best trade-off between counting accuracy and navigation cost.
+Code can be found at https://github.com/longrunling/ECC?.
 1 Introduction
-Crowd counting is critical for public safety and urban planning [21].
+Crowd counting is critical for public safety and urban planning kang2018beyond .
 One main challenge in this field is occlusion.
@@
 To summarize, these situations are caused by a poor observation point.
-Existing methods try to solve these challenges from several perspectives, such as using multi-scale feature extraction [20], body part detection [26] or multi-camera fusion [49, 50].
-Datasets are also expanded for enhancing method performance [51, 19, 36, 43].
-In general, these methods or datasets either try to enhance learned model representation, or expand the camera sensing range by introducing multi-view settings.
+Existing methods try to solve these challenges from several perspectives, such as using multi-scale feature extraction kang2018crowd , body part detection liu2019point , or multi-camera fusion zhang2019wide ; zhang2024multi .
+Datasets are also expanded for enhancing method performance zhang2016single ; idrees2018composition ; sindagi2020jhu ; wang2020nwpu .
+In general, these methods or datasets either try to enhance learned model representation or expand the camera sensing range by introducing multi-view settings.
 Yet, their passive camera settings do not fully solve the occlusion challenge in crowd counting, especially in cases where crowds exceed the sensing range, or no cameras are set to detect completely obscured crowds.
 These issues restrict passive-camera-based methods in practice.
-Recent development in Embodied AI brings perspectives to solve the occlusion challenge in crowd counting.
+Figure 1: (a) Comparison between ECCD and existing embodied navigation datasets.
+ECCD features large-scale outdoor crowd scenes.
+(b) Comparison between ECCD and crowd counting datasets.
+ECCD enables interactive ability.
+(c) Comparison between ZECC and existing crowd counting methods.
+ZECC is an agentic framework with automatic camera adjusting ability.
+Recent development in Embodied AI brings a new perspective to address the occlusion challenge in crowd counting.
 It has been demonstrated to possess significant potential in enhancing scene exploration and object detection.
-Aspects such as Vision-Language-Navigation (VLN) have been directed towards equipping mobile robots with human-like perception abilities, resulting in remarkable performance in exploring scenes and detecting objects in open environments [9, 4, 52, 7].
-The active camera settings in VLN is promising to solve the fundamental challenges in crowd counting, since it can optimize observation point to mitigate overlap and invisibility caused by poor camera position setting.
-Yet most VLN benchmarks [44, 40, 3, 1] are indoor environments with limited exploration space (e.g., no Z-axis action options) and relatively small object quantity.
+Aspects such as Vision-Language-Navigation (VLN) have been directed towards equipping mobile robots with human-like perception abilities, resulting in remarkable performance in exploring scenes and detecting objects in open environments gadre2023cows ; chen2023typefly ; zhao2023agent ; choi2024find .
+The active camera settings in VLN are promising to solve the fundamental challenges in crowd counting, since it can optimize observation points to mitigate overlap and invisibility caused by fixed camera position settings.
+Yet most VLN benchmarks wang2024embodiedscan ; wald2019rio ; chang2017matterport3d ; brazil2023omni3d are indoor environments with limited exploration space (e.g., no Z-axis action options) and relatively small object quantity.
 And the performance of such methods remains unknown for detecting crowds with a large quantity and complex distribution in large-scale scenes.
 This results in a significant gap between VLN and crowd counting, as in practice, crowds often appear in large spaces with varying distribution.
-To address the issues mentioned above, we first define a novel task, Embodied Crowd Counting (ECC).
-This task integrates crowd counting into the Zero-shot Object-Goal Navigation task (ZSON) in VLN [28].
-It is designed for large-scale scenes with a large object quantity.
-A testing environment, Embodied Crowd Counting Dataset (ECCD), is then developed for this task.
-Compared to existing crowd counting and VLN datasets, this test environment has the following features: 1) It supports 60 distinct and diverse large-scale outdoor scenes.
-Each scene covers an area up to 40,000 $m^{2}$, with target quantity up to 15,000.
-2) It uses a prior distribution, Poisson Point Process [11], for crowd quantity distribution modeling, which approximates realistic crowd situations.
-Based on this platform, we develop a method, Zero-shot Embodied Crowd Counting (ZECC), to conduct crowd counting in ECCD.
-Compared to existing crowd counting methods and VLN methods, ZECC has the following features: 1) It uses active camera setting to solve fundamental challenges in crowd counting, transferring the detection task to an exploration task; 2) It consists of a coarse-to-fine navigation mechanism which uses common sense of Multi-Modal-Large-Language-Model (MLLM) to analyze the surrounding environment and plan Z-axis exploration, leveraging advantages of 6 degrees of freedom (6-DoF) to avoid low-altitude obstacles and gain border sight, thus achieving efficient exploration; 3) It consists of a crowd distribution analysis mechanism which transfers the complex crowd distribution to simple surfaces.
-The normal lines of the surfaces are used to generate fine observation points, which are used to mitigate the influence of overlap and invisibility, enabling accurate crowd counting.
+To address the issues, we first define a novel task, Embodied Crowd Counting (ECC), which is shown in Figure 1.
+The task is defined as counting the total number of people present in a large outdoor scene using a drone.
+Given the absence of an existing dataset, we have created a new dataset called the Embodied Crowd Counting Dataset (ECCD) specifically for this task.
+This dataset includes 60 unique and diverse large-scale outdoor scenes.
+Each scene spans an area of up to 40,000 $m^{2}$ and has a target crowd size of up to 15,000 individuals.
+To ensure realism in our dataset, we employ a Poisson Point Process ge2009marked for modeling the distribution of crowd sizes, which effectively simulates real-world crowd scenarios.
+The comparision of ECCD and related datasets are shown in Table 1.
+Table 1: Comparison between ECCD and other related datasets. ECCD combines features from both crowd counting datasets and embodied navigation datasets.
+Dataset Active camera Place Max target quantity / sample Dynamic target w/o Instruction Task
+NWPU-Crowd wang2020nwpu $\times$ - 20,033 $\times$ - $C$
+SenseCrowd li2022video $\times$ - 296 $\checkmark$ - $C$
+DukeMTMC ristani2016performance $\times$ - 2,834 $\checkmark$ - $C$
+R2R anderson2018vision $\checkmark$ Indoor (Ground) - $\times$ $\times$ VLN
+HM3DSem ramakrishnan2021habitat $\checkmark$ Indoor (Ground) - $\times$ $\checkmark$ ObjNav
+AerialVLN gao2024aerial $\checkmark$ City(Aerial) - $\times$ $\times$ VLN
+CityNav lee2024citynav $\checkmark$ City(Aerial) - $\times$ $\times$ VLN
+Openfly gao2025openfly $\checkmark$ City(Aerial) - $\times$ $\times$ VLN
+ECCD $\checkmark$ City(Aerial) 15,488 $\times$ $\checkmark$ ECC
+In this study, we introduce a baseline method, Zero-shot Embodied Crowd Counting (ZECC) aiming at counting individuals present in environments populated by crowds.
+Given the ability of interacting dynamically with surroundings using foundation models, modern argentic methods achieve even better performance against training-based methods while maintaining generalization ability.
+gadre2023cows ; kuang2024openfmnav ; gao2024aerial ; long2024instructnav .
+Inspired by this, we aim to build a zero-shot baseline that can generate to diverse scenes leveraging foundation models, using an argentic paradigm without a vast amount of training data.
+The primary challenge lies in choosing suitable navigation points to find a balance between efficient exploration and effective detection.
+To address this, our approach consists of two main components: the Active Top-down Exploration (ATE) method and the Normal-line Based Navigation (NLBN) system.
+The ATE method serves as an efficient exploration strategy that employs a coarse-to-fine navigation approach.
+It utilizes the common sense capabilities of Multi-Modal Large Language Models (MLLM) to assess the environment, allowing for effective planning of vertical movement (Z-axis exploration).
+This takes advantage of the six degrees of freedom (6-DoF), which helps to avoid obstacles at lower altitudes and provides a broader view of the surroundings, thereby facilitating better exploration.
+Following the rough estimation of crowd distribution produced by ATE, we propose the NLBN to create precise navigation points for improved crowd detection.
+By using the normal lines of surfaces, we can establish detailed observation points that achieve a balanced trade-off between exploration efficiency and detection performance.
+This method addresses challenges related to overlapping individuals and visibility issues, ultimately leading to more accurate crowd counting.
+Experimental results show that the method is effective due to its interactive ability.
 The contributions of this work can be summarized as follows:
@@
 Unlike traditional crowd counting and VLN datasets, it features large-scale outdoor crowd scenes with interactive capabilities.
-  • ZECC is proposed to use MLLM for Z-axis exploration, reducing costs while ensuring detection performance.
-By utilizing normal lines to calculate navigation points, this approach eliminates overlap and enhances visibility in crowds.
-  • Experiments show that the proposed method outperforms the baseline methods in various testing environments.
-2 Related Work
+  • We propose a baseline method, ZECC, using MLLM for Z-axis exploration, reducing costs while ensuring detection performance.
+By utilizing normal lines to calculate navigation points, this approach eliminates occlusion and enhances visibility in crowds.
+2 Related works
 2.1 Crowd Counting
-Crowd-counting algorithms have greatly benefited from large-scale, high-quality datasets like UCF-CC50 [18] and UCF-QNRF [19].
-These foundational datasets have facilitated the creation of subsequent collections focused on dense crowd imagery, such as ShanghaiTech [51], JHU-CROWD++ [36], NWPU-Crowd [43].
+Crowd-counting algorithms have greatly benefited from large-scale, high-quality datasets like UCF-CC50 idrees2013multi and UCF-QNRF idrees2018composition .
+These foundational datasets have facilitated the creation of subsequent collections focused on dense crowd imagery, such as ShanghaiTech zhang2016single , JHU-CROWD++ sindagi2020jhu , NWPU-Crowd wang2020nwpu .
 However, the images in these datasets are generated from fixed cameras.
-In contrast, ECCD provides interactive capabilities while maintains diverse crowd distribution.
-Methods like [51, 20, 24, 24, 5] leverage crowd distribution prior in an image, or using attention maps to learn dependency.
-Recent multi-modal approaches [10, 47, 29, 42] leverage vision-language models to transfer image-text knowledge to dense crowd prediction.
-While these models improve long range and overlapping small target detection ability, their performance are restricted in practical if overlap reaches extreme level.
-Recent efforts expand spatial coverage.
-These works include multi-view systems [32, 17] that use multi camera to capture images from larger scale scenes.
-Others like [14, 15] use recorded video to conduct crowd analysis.
-While these are important advances on solving the fundamental challenges in crowd counting against the single-fix-image-based methods, their camera setting is still predefined by the dataset and cannot be adjusted when inferring.
-This cannot ensure fully exploration in a large environment.
+In contrast, ECCD provides interactive capabilities while maintaining diverse crowd distribution.
+Methods like zhang2016single ; kang2018crowd ; liu2023point ; chen2024improving leverage crowd distribution prior in an image, or use attention maps to learn dependency.
+Recent multi-modal approaches gao2024clip ; yao2024cpt ; meng2024multi ; wang2024multi leverage vision-language models to transfer image-text knowledge to dense crowd prediction.
+While these models improve long-range and overlapping small target detection ability, their performance is restricted if the overlap reaches an extreme level.
+Recent efforts expand spatial coverage by multi-view systems qiu20223d ; hu2024mvd that use multiple cameras to capture images from large-scale scenes.
+Others like han2022dr ; han2023change use recorded video to conduct crowd analysis.
+Although these advances represent significant progress in addressing the basic challenges of crowd counting compared to traditional methods that rely on fixed images, the camera settings remain predetermined by the dataset.
+This restriction means that the settings cannot be modified during the inference process, limiting the ability to fully examine larger environments.
 ZECC allows active exploration, which is fundamentally different from existing methods.
 2.2 Embodied Navigation
-Many traditional robot navigation methods were developed using conventional navigation datasets like KITTI [12] and SUN RGB-D [38].
-Beyond these foundational datasets, [44, 33] provide 3D indoor environments for navigation and interaction tasks.
+Many traditional robot navigation methods were developed using conventional navigation datasets like KITTI geiger2012we and SUN RGB-D song2015robot .
+Beyond these foundational datasets, wang2024embodiedscan ; ramakrishnan2021habitat provide 3D indoor environments for navigation and interaction tasks.
 These datasets are limited with their small scale size and small object quantity.
-Recent datasets have been created towards outdoor navigation like AerialVLN [25], and CityNav [23].
-However, they are designed for VLN tasks without consideration of object quantity.
-Compared to these datasets, ECCD simultaneously supports large scale outdoor scenes and large object quantity with diverse distribution.
+Recent datasets have been created towards outdoor navigation liu2023aerialvln ; lee2024citynav ; gao2025openfly ; liu2025aerialvg However, they are designed for VLN tasks without consideration of object quantity.
+Compared to these datasets, ECCD simultaneously supports large-scale outdoor scenes and large object quantity with diverse distribution.
 Efficient exploration using mobile robots remains a crucial challenge in vision and robotics.
-[13, 16, 30, 38] have developed human-like cognitive maps, enabling autonomous path learning in unknown environments.
-Other approaches like [2, 6, 31] use reinforcement learning to develop exploration policies.
-Recently, [48] applied video-based visual-language models to plan sequential actions in VLN.
-And [27, 53] presented zero-shot models that utilizes natural language instructions to guide agents through environments without prior environment-specific training.
-However, the methods mentioned above are for indoors navigation without Z-axis moving ability.
-This restricts their ability in large scale outdoor scenes.
-Recently, methods like [25, 45] introduce 6-DoF in outdoor VLN.
-Yet they are designed for instruction following task, in which the movement is restricted by human language.
-Besides, they lack ability to analyze crowds with large quantity and complex distribution.
-In contrast, ZECC is the first self-motivated agent in 6-DoF and can handle crowds simultaneously.
-Figure 2: The proposed framework for ZECC consists of several key steps.
-First, we utilize an exploration strategy based on FBE to create a global crowd distribution by generating a target point cloud.
-Next, we apply a Gaussian Mixture Model (GMM) clustering to identify the representative centers of the global target point cloud and determine the final navigation points based on potential navigation directions.
-Finally, we integrate detection results obtained from various detectors with depth observations, mapping these to the global point cloud.
-The number of filtered targets is then counted and recorded as the final result for the crowd.
+gupta2017cognitive ; henry2012rgb ; newcombe2011kinectfusion ; song2015robot ; gao2024aerial have developed human-like cognitive maps, enabling autonomous path learning in unknown environments.
+Other approaches like burda2018exploration ; chen2019learning ; parisi2021interesting use reinforcement learning to develop exploration policies.
+Recently, zhang2024navid applied video-based visual-language models to plan sequential actions in VLN.
+And long2024instructnav ; zhou2024navgpt presented zero-shot models that utilize natural language instructions to guide agents through environments without prior environment-specific training.
+However, the methods mentioned above are for indoor navigation without Z-axis moving ability.
+This restricts their ability in large-scale outdoor scenes.
+Recently, methods like liu2023aerialvln ; wang2024towards introduce 6-DoF in outdoor VLN.
+Yet they are designed for instruction following tasks, in which the movement is restricted by human language.
+Besides, they lack the ability to analyze crowds with a large quantity and complex distribution.
+In contrast, ZECC is the first self-motivated agent in 6-DoF that can handle crowds.
 3 Method
@@
 The counting error and travel distance are calculated to assess the agent’s performance.
-ECC can be considered analogous to the ZSON task, since they both require an agent to detect targets in an unexplored environment without additional assistance.
-However, there are several differences.
-The task setting differences are: 1) The object category $G$ in ZSON is limited to crowd in ECC.
-2) The step limitation in ZSON is removed in ECC.
-3) The evaluation criterion in ECC is changed to counting error and travel distance.
-These differences do not prevent the availability of most recent ZSON methods in ECC by conducting minor modifications.
-However, agents should face essential differences in ECC: 1) Z-axis is available, which is not considered by most recent ZSON methods.
-2) Complex crowd distributions exist, including heavy occlusions, while ZSON methods only consider objects with small quantity.
+ECC can be considered analogous to the Zero-Shot Object-Goal Navigation (ZSON) task majumdar2022zson , since they both require an agent to detect targets in an unexplored environment without additional assistance.
+However, ECC faces unique challenges: 1) The Z-axis is available, which is not considered by most ZSON methods.
+2) Complex crowd distributions exist, including heavy occlusions, while ZSON methods only consider fewer objects.
 Current ZSON methods show unknown performance under such differences, and new methods should be considered in ECC.
-Note that ECC is different from the instruction following tasks in VLN such as [25, 45], since these tasks require natural language assistance to drive the agent.
+Note that ECC is different from the instruction following tasks in VLN such as liu2023aerialvln ; wang2024towards , since these tasks require natural language assistance to drive the agent.
 3.2 Embodied Crowd Counting Dataset (ECCD)
-We propose a new dataset called ECCD to support algorithm design and evaluation for the community.
-ECCD is developed using Unreal Engine 4.
+We propose a new dataset called ECCD to support algorithm design and evaluation for the community, developed using Unreal Engine 4.
 This platform enables programmable environment construction, allowing scene richness and scalability.
@@
 Also, ECCD has an area up to 40,000 $m^{2}$, reaches a height up to 50 $m$, and allows for the simultaneous presence of over 15,000 targets within a single scene.
-This is significantly different from existing robot navigation datasets, such as [33] which offers environments with an average navigable space of 1,000 $m^{2}$, and crowd counting datasets, such as [51] which contains 1,198 annotated images captured with static cameras and a total of 330,165 individuals.
+This is significantly different from existing robot navigation datasets, such as ramakrishnan2021habitat , which offers environments with an average navigable space of 1,000 $m^{2}$, and crowd counting datasets, such as zhang2016single , which contains 1,198 annotated images captured with static cameras and a total of 330,165 individuals.
 Realism.
-The ECCD is designed to simulate large-scale outdoor crowd scenarios in reality, in order to ensure practical effectiveness of systems built on ECCD, as shown in Figure 3.
+The ECCD is designed to simulate large-scale outdoor crowd scenarios in reality, in order to ensure practical effectiveness of systems built on ECCD, as shown in Figure 2 (a).
 Environments like a city, a stadium, and a parking lot are simulated.
@@
 These features allow ECCD to reflect challenges in reality.
-Figure 3: ECCD is designed to realistically mimic building and crowd distribution. On the left are samples from ECCD, and on the right are the real scenes.
+Figure 2: (a) ECCD is designed to mimic building and crowd distribution realistically.
+On the left are samples from ECCD, and on the right are the real scenes.
+(b) Illustration of the potential navigation vectors, normal lines, and FBE view vectors.
+Zoom in for better visualization.
 Crowd generation mechanism.
-To model real challenges in crowd counting, ECCD uses Poisson Point Process for crowd quantity distribution modeling [11].
-In ECCD, Blocks are placed in the simulator to present potential areas where crowds may exist.
+To model real distribution in crowd counting, ECCD uses Poisson Point Process for crowd quantity distribution modeling ge2009marked .
+In ECCD, Blocks are placed in potential areas where crowds may exist.
 Then, for each block $\mathcal{U}\subset\mathbb{R}^{2}$, the process is defined as:
@@
 3.3 Zero-shot Embodied Crowd Counting (ZECC)
+Figure 3: The proposed framework.
+First, ATE is proposed to estimate the global crowd distribution efficiently.
+Then, NLBN is proposed to generate fine observation points, alleviating crowd overlap.
+The final result is generated by aggregating all fine detections.
 3.3.1 Overview
-Previous embodied navigation agents are designed for indoor environments [27], or relying on language assistance [25, 45], making them restricted in large-scale outdoor scenes.
-Under such context, a zero-shot agent, ZECC, which can actively control altitude and conduct crowd analysis, is proposed.
-As illustrated in Figure 2, the method consists of 2 components, Active Top-down Exploration (ATE) and Normal-line based Navigation (NLBN).
-ATE is designed for adjusting agent altitude for efficient coarse crowd distribution estimation, and NLBN estimates normal lines on top of crowd for accurate crowd observation, alleviating occlusion.
+Previous embodied navigation agents are designed for indoor environments long2024instructnav , or relying on language assistance liu2023aerialvln ; wang2024towards , making them restricted in large-scale outdoor scenes.
+Under such a context, a zero-shot agent, ZECC, which can actively control altitude and conduct crowd analysis, is proposed.
+As illustrated in Figure 3, the method consists of three components: Active Top-down Exploration (ATE), Normal-line based Navigation (NLBN), and Fine detection.
+ATE is designed for adjusting agent altitude for efficient coarse crowd distribution estimation, and NLBN estimates normal lines on top of the crowd for accurate crowd observation, alleviating occlusion.
 3.3.2 Active Top-down Exploration (ATE)
 Crowds are often concentrated in specific areas, such as roads and squares, making it unnecessary to explore every region of the environment.
-To address this, ATE is proposed to estimate the global distribution of crowds from a high altitude.
-This approach aims to improve efficiency by focusing exploration efforts on the most relevant areas and avoiding unnecessary locations.
-In particular.
-high-altitude exploration (HAE) brings a broader field of view for decision making, as well as relatively less exploration cost since obstacles are often sparse in high altitude.
-While low-altitude exploration (LAE) provides close range observation for clear target detection.
-It is essential for the agent to plan HAE and LAE to achieve efficiency and accuracy simultaneously.
-Therefore, ATE leverages the Z-axis mobility of outdoor agents and the common sense of multi-modal large language model (MLLM) to switch between HAE and LAE by local environment layout reasoning.
-Then, the crowd distribution is estimated by crowd counting model to predict density maps in the environment.
+To address this, ATE is proposed to estimate global crowd distribution by changing altitude.
+This approach aims to improve efficiency by focusing exploration efforts on the most relevant areas.
+In particular, high-altitude exploration (HAE) brings a broader field of view for decision making, as well as relatively less exploration cost since obstacles are often sparse in high altitude.
+In contrast, low-altitude exploration (LAE) provides close-range observation for precise target detection.
+The agent needs to plan HAE and LAE to achieve efficiency and accuracy simultaneously.
+Therefore, ATE leverages the Z-axis mobility of outdoor agents and the common sense of MLLM to switch between HAE and LAE using local environment layout reasoning.
+Then, the crowd distribution is estimated by a crowd counting model to predict density maps.
 Specifically, the agent collects observations ${O_{t}}=\left\{{o_{t}^{1},\ldots,o_{t}^{c}}\right\}$ and pose $p_{t}$ at time step $t$, where $c$ indicates the camera number of the agent.
-Then, a MLLM is prompted by using the observations and text prompts to conduct environment layout reasoning.
-The MLLM is asked to predict whether current location is valuable for LAE, by refering current crowd appearence and obstacle layout.
-This process is formulate as:
+Then, an MLLM is prompted using observations and text prompts to conduct environment layout reasoning.
+The MLLM is asked to predict whether the current location is valuable for LAE, by referring to the current crowd appearance and obstacle layout.
+This process is formulated as:
 $s_{t}=\text{MLLM}\left({{O_{t}};I}\right),$ (2)
@@
 If $s_{t}>0.5$, the agent will adjust its altitude for LAE.
-After switching altitude strategy, the agent will conduct regular exploration.
-For HAE, one of the frontiers between explored and unknown areas is selected as navigation point.
-For LAE, the agent will keep exploring frontiers until areas within its field of view during HAE are fully explored.
-During LAE, once the agent reaches a frontier $f$, it gathers observations ${O_{f}}=\left\{{o_{f}^{1},\ldots,o_{f}^{c}}\right\}$, and a crowd counting model predicts crowd density maps on the observations.
-Then the density maps are projected onto the global point cloud to form a global crowd distribution.
-This is formulated as:
+After switching the altitude strategy, the agent will conduct regular exploration.
+For HAE, one of the frontiers between explored and unknown areas is selected as a navigation point.
+For LAE, the agent keeps exploring until areas within its field of view during HAE are fully explored.
+During LAE, once the agent reaches a navigation site $f$, it gathers observations ${O_{f}}=\left\{{o_{f}^{1},\ldots,o_{f}^{c}}\right\}$, and a crowd counting model predicts crowd density maps on the observations.
+Then, the density maps are projected onto the global point cloud to form a global crowd distribution:
 $d_{f}=\text{P}\left({\text{G}\left({{O_{f}}}\right),{p_{f}}}\right),$ (3)
@@
 3.3.3 Normal-line based Navigation (NLBN)
-NLBN is utilized for the precise analysis of potential overlapping structures within crowds.
-The intuition behind this approach is that while a random viewpoint may often obscure visibility in a dense crowd, there are several vantage points that can still provide a clear view for identifying individuals.
-These observation points are expected to be above and maintain a certain degree to the center of the crowd, since a top-down view can distinguish overlapping, and perspectives from certain angles can ensure that the targets are in the field of view (FOV) of the agent.
-To find such observation points, NLBN transfers the crowd detection task into a surface detection task.
-By fitting surfaces on the crowd distribution, normal lines of the surfaces can be obtained, and view points which fulfill the constraints mentioned above can be accurately calculated by using normal lines.
-After achieving global crowd distribution in ATE, NLBN first clusters the crowd point cloud into subregions using a clustering method.
-This step helps divide the large crowd point cloud into manageable parts, as navigating to each navigation point can be resource-intensive.
-Next, surfaces are fitted to these subregions based on the point cloud data to obtain normal lines.
-This process transforms the challenging task of crowd analysis into a more manageable surface analysis.
-Finally, optimized navigation points are generated based on the normal lines, ensuring they maintain a certain distance and angle from the crowd surfaces for accurate crowd detection.
-A viewpoint-based method is employed to confirm that the selected navigation points do not overlap with any obstacles.
-In particular, GMM [35] is used to divide the global crowd distribution into manageable subregions.
-It ensures each sample point in a distribution can be assigned with a cluster.
-Thus, it can divide any crowd distribution into patches.
+NLBN is designed to analyze overlapping structures in dense crowds by converting the crowd detection task into surface detection, which helps identify optimal observation points.
+While random viewpoints may impair visibility, vantage points located above the center of the crowd provide clearer views for individual identification.
+This top-down perspective enables the distinction of overlapping individuals and maintains targets within the field of view (FOV).
+Initially, NLBN clusters the crowd point cloud into subregions using a clustering method, simplifying the analysis by breaking the large point cloud into manageable parts, as navigating to each point can be resource-intensive.
+Surfaces are then fitted to derive normal lines, thus transforming complex crowd analysis into a more straightforward surface analysis.
+Finally, optimized navigation points are generated based on these normal lines, ensuring they are elevated to facilitate accurate crowd detection while employing a viewpoint-based approach to avoid overlaps with obstacles.
+In particular, GMM scrucca2023introduction is used to divide the global crowd distribution into manageable subregions.
+It can divide any crowd distribution into patches.
 A parameter $\epsilon$ is used to determine the size of each GMM cluster.
-After the clusters and centers are generated, surfaces are fitted for each patch.
+Surfaces are fitted for each patch afterwards.
 Then, for the $N$-th cluster, the normal is obtained and represented by ${\bf{d}}_{N}^{{\rm{cluster}}}$.
@@
 where $\zeta$ is a hyper parameter.
-These candidate view directions can bring the agent to a optimized observation point by selecting a position along the vector.
+These candidate view directions can bring the agent to optimized observation points by selecting a position along the vector.
 However, there are two issues: 1) The position may be located on obstacles; 2) The crowd cluster may be out of the agent’s FOV.
-To solve the issues and identify the final navigation points from the potential ones, a simple yet effective technique is proposed.
-In the process of ATE, each cluster point identified by the counting model processes at least one viewpoint of the agent.
-The vector connecting the viewpoints and target points firmly indicates navigable space, since the propagation of light naturally points to unobstructed area.
-Besides, the vector can ensure that the crowd is in the FOV as well, since the target image is identified during ATE.
-Based on this prior, the ATE viewpoints are used to generate the final navigation points.
-The ATE view vector is calculated as:
+Since the propagation of light naturally points to unobstructed areas, the ATE viewpoints are used to generate the final navigation points, which are calculated as:
 $\mathbf{\bf{d}}_{N}^{{\rm{ATE}}}={\bf{x}}_{N}^{{\rm{ATE}}}-{\bf{x}}_{N}^{{\rm{cluster}}},$ (5)
@@
 NLBN ensures close-range, precise target observation and safe navigation, even in complex and unfamiliar environments.
-The potential navigation vectors, normal lines, and ATE view vectors are illustrated in Figure 4.
-Figure 4: Illustration of the potential navigation vectors, normal lines, and FBE view vectors. Zoom in for better visualization.
+The potential navigation vectors, normal lines, and ATE view vectors are illustrated in Figure 2 (b).
 3.3.4 Fine Detection and Counting (FDC)
 Using the navigation points generated by NBLN, the agent travels from one point to another through path planning algorithms.
-Upon reaching each navigation point, close-range and high-resolution RGB observations can be conducted.
-These observations are then utilized as inputs for detection models to perform precise target detection.
-The detection results are projected onto the global point cloud using the depth sensor and the agent pose.
+Close-range and high-resolution RGB observations can be conducted upon reaching each navigation point.
+These observations are then utilized for detection models to perform precise target detection.
+The detection results are projected onto the global point cloud using the depth sensor and the agent’s pose.
 To prevent repetitive detections, only one target is retained for each region within a specific scale.
-Ultimately, the result is calculated by counting the quantity of filtered targets.
+Ultimately, the result is calculated by counting the number of filtered targets.
 4 Experiments
-4.1 Settings
-Metrics. Mean Absolute Percentage Error (MAPE) is used as the metric to evaluate the counting performance, which is defined as:
-[…]
-where $M$ is the quantity of testing environments, $\hat{y}_{i}$ and $y_{i}$ are the estimated count and the ground truth count in an environment, respectively.
-The sum of Euclidean distance between adjacent navigation points along the agent traveling path is used to evaluate the travel distance TD, which is defined as:
-${\rm{TD}}=\sum\limits_{i=1}^{n-1}{\left\|{{{\bf{x}}_{i+1}}-{{\bf{x}}_{i}}}\right\|},$ (9)
-where $n$ is the quantity of navigation points in one episode, ${{{\bf{x}}_{i}}}$ and ${{{\bf{x}}_{i+1}}}$ are the coordinates of the two adjacent navigation points, respectively.
+Baselines.
+We compare ZECC with exploration methods: Frontier-based exploration (FBE) topiwala2018frontier , ZSON methods: CoW gadre2023cows and OpenFMNav kuang2024openfmnav , and multi-view counting (MVC) methods zhang2019wide ; mo2024countformer .
+Metrics.
+Mean Absolute Percentage Error (MAPE) is used to evaluate the counting performance:  $\text{ MAPE }=\frac{1}{M}\sum_{i=1}^{M}\left|\frac{y_{i}-\hat{y}_{i}}{y_{i}}\right|\times 100\%$ where $M$ is the quantity of testing environments, $\hat{y}_{i}$ and $y_{i}$ are the estimated count and the ground truth count, respectively.
+For ZSON methods, the sum of Euclidean distance between adjacent navigation points along the agent traveling path is used to evaluate the travel distance (TD), which is defined as: ${\rm{TD}}=\sum\limits_{i=1}^{n-1}{\left\|{{{\bf{x}}_{i+1}}-{{\bf{x}}_{i}}}\right\|},$ where $n$ is the quantity of navigation points in one episode, ${\bf{x}}_{i}$}} and ${\bf{x}}_{i+1}$}} are the coordinates of the two adjacent navigation points, respectively.
+For the multi-view crowd counting method, we report the number of cameras required to achieve comparable performance to the proposed approach.
 Implementation Details.
-MLLM used in ATE is GPT-4V [46].
-The path planning method is A* algorithm [37].
+MLLM used in ATE is GPT-4V yang2023dawn .
+smolka2019pathfinding is used as path planning method.
 Altitude is 80m for HAE and 10m for LAE.
-The crowd density estimator in ATE is Generalized Loss (GL) [41], and the detection model in FDC is Grounding DINO (GD) [34].
+The crowd density estimator in ATE is Generalized Loss (GL) wan2021generalized , and the detection model in FDC is Grounding DINO (GD) ren2024grounding .
 For hyper parameters, navigation vector deg $\zeta$ is 15 °, density threshold $\kappa$ is 0.7, navigation point range $\eta$ is 8 m, and cluster size $\epsilon$ is 40.
-Baselines.
-We compare ZECC with existing exploration method Frontier-based exploration (FBE) [39], as well as SOTA ZSON methods, 1) CoW [8], 2) OpenFMNav [22].
-These methods use advanced foundation models to conduct decision making during navigation.
-To apply the baselines in ECC, the exploration step limitation for ZSON methods is removed, and they are equipped with GD or GL to conduct crowd counting using the projection method in FDC.
-4.2 Results
-Overall Performance.
-We report overall performance in Table 1.
+For the ZSON methods, the exploration step limitation is removed.
+When the methods finish an exploration step, they additionally get RGB-D captures of the environment.
+As exploration stops, they are equipped with GD or GL to conduct crowd counting using the projection method in FDC upon their captured images.
+For MVC, the scenes are divided into grids with diverse intervals, which are 10m, 20m, and 30m.
+Four cameras at poses of 0°, 90°, 180°, and 270° are placed on each grid intersection to obtain RGB captures.
+Then the captured images are sent to the MVC to get counting results.
+All methods are proceeded on a platform with IntelCorei9-14900KF, 128GBRAM, and NVIDIA GeForce RTX 4090 GPU.
+5 Results
+5.1 Overall Performance
+Comparison with ZSON methods.
+We report performance in Table 3.
 ZECC offers the optimal balance between counting performance and cost.
-In contrast, FBE delivers the best TD but lacks the ability to perform a target detection function, which limits its capacity for fine observation.
-This incomplete observation negatively impacts its overall performance.
-ZSON methods, such as Cow and OpenFMNav come with perception modules; however, they do not select a 3D fine observation point as NLBN, which means they cannot fully observe each individual in the crowd.
-This limitation leads to individuals being obscured by overlapping crowds, especially when the density is high.
-In some instances, these methods only identify single individuals, which reduces their efficiency.
-In contrast, ZECC improves counting efficiency by planning Z-axis exploration using ATE and employing NLBN for detailed detection.
-For agents utilizing counting models for crowd counting, the performance tends to be slightly lower than that of agents using detection models.
+In contrast, FBE delivers the best TD but cannot perform a target detection function, which limits its capacity for fine observation.
+ZSON methods come with perception modules; however, they do not select a 3D fine observation point as NLBN, which means they cannot fully observe each individual in the crowd.
+Agents utilizing counting models for crowd counting tend to perform slightly lower than agents using detection models.
 This is due to the increased sparsity of crowd observations as the agents get closer to the crowd.
 Crowd counting models are mainly trained on images depicting dense crowds, which results in limited generalization capability for scenarios with sparser crowds.
-Additionally, current ZSON methods do not consider finding a close position for crowd observation, which further degrades performance.
+Table 2: Comparison with ZSON methods. ZECC achieves a trade-off between MAPE and TD.
 Method MAPE (%) TD (m)
-FBE [39] + GL [41] 55.92 2227.71
-FBE [39] + GD [34] 53.85 2227.71
-CoW [8] + GL [41] 51.19 3271.07
-CoW [8] + GD [34] 46.70 3271.07
-OpenFMNav [22] + GL [41] 59.71 4936.79
-OpenFMNav [22] + GD [34] 46.84 4936.79
-ZECC 18.91 3804.63
-Table 1: Overall performance of the methods. ZECC achieves best trade-off between MAPE and TD.
-Figure 5: Performance and cost of ZECC and the baselines under different crowd density levels.
-L1-L5 refers to increasing density level.
-The figure demonstrate that ZECC achieves a balance between performance and exploration cost.
-Figure 6: The effect of four hyperparameters used in our model, including GMM cluster size, navigation vector degree, navigation point range and density map threshold.
+FBE topiwala2018frontier + GL wan2021generalized 57.19 ± 1.83 2513.06 ± 247.35
+FBE topiwala2018frontier + GD ren2024grounding 53.38 ± 1.26 2513.06 ± 247.35
+CoW gadre2023cows + GL wan2021generalized 52.75 ± 1.52 3449.51 ± 127.2
+CoW gadre2023cows + GD ren2024grounding 46.01 ± 0.96 3449.51 ± 127.2
+OpenFMNav kuang2024openfmnav + GL wan2021generalized 60.57 ± 2.43 5069.64 ± 183.23
+OpenFMNav kuang2024openfmnav + GD ren2024grounding 49.41 ± 2.35 5069.64 ± 183.23
+ZECC 18.71 ± 1.41 3722.45 ± 73.78
+Table 3: Comparison with MVC methods. ZECC achieves a trade-off between MAPE and cost.
+Method MAPE (%) # of Cameras
+MVF-10 zhang2019wide 15.13 ± 0.00 1735.32 ± 0.00
+MVF-20 zhang2019wide 39.92 ± 0.00 747.32 ± 0.00
+MVF-30 zhang2019wide 61.43 ± 0.00 333.32 ± 0.00
+CountFormer-10 mo2024countformer 12.8 ± 0.00 1735.32 ± 0.00
+CountFormer-20 mo2024countformer 35.26 ± 0.00 747.32 ± 0.00
+CountFormer-30 mo2024countformer 56.76 ± 0.00 333.32 ± 0.00
+ZECC 18.71 ± 1.41 5 ± 0.00
 We analyze various environments with different crowd density levels and visualize the average performance and cost of the methods using the GD detector in these scenarios.
@@
 Among the methods evaluated, ZECC demonstrates the best average performance across different density levels.
-As crowd density increases, the MLLM in ATE becomes more effective at identifying high-density crowds, resulting in improved performance.
-In contrast, other methods struggle to manage high-density crowds, leading to limited performance.
+As crowd density increases, the MLLM in ATE with NLBN becomes more effective at identifying and observing high-density crowds, resulting in improved performance.
+In contrast, other methods struggle to detect high-density crowds.
+Figure 4: Performance and cost of ZECC and the baselines under different crowd density levels.
+L1-L5 refers to increasing density level.
+The figure demonstrates that ZECC achieves a balance between performance and exploration cost.
 In terms of cost, ZECC’s navigation points are influenced by crowd distribution and density.
@@
 Although ZECC falls short of achieving the lowest cost in the last two density levels, its costs are still comparable to the baseline while ensuring effective counting performance.
-In contrast, other methods do not actively adjust navigation points based on crowd distribution, which results in their costs remaining relatively stable and limits their effectiveness in low-density scenes.
-MAPE (%) TD (m)
-ZECC w/o ATE 17.46 4633.67
-ZECC w/o HAE 17.46 4633.67
-ZECC w/o LAE 88.08 1738.84
-ZECC 18.91 3804.63
-Table 2: The results of ablation study for components in ATE.
-w/o ATE refers to using FBE + NLBN results for crowd counting.
-w/o HAE refers to fix the agent’s altitude to LAE.
+In contrast, other methods do not actively adjust navigation points.
+Their costs remain relatively stable, yet their performance is limited.
+Comparison with MVC.
+The comparison with MVC is shown in Table 3.
+"-10" refers to MVC methoods using a grid interval of 10m.
+CountFormer-10 achieves the best performance, while ZECC still provides a close performance by reducing the camera used significantly.
+This illustrates the advantage of the active method over the multi-view method.
+5.2 Ablation Study
+ATE.
+We conducted ablation studies by removing specific components from ATE.
+The results are presented in Table 5.
+w/o HTE refers to using FBE + NLBN results for crowd counting.
 w/o LAE refers to fix the agent’s altitude to HAE.
-ZECC improves approximately 17% cost by reducing approximately 8% performance, achieving a better trade-off.
-MAPE (%) Successful rate (%)
-ZECC w/o NLBN 65.19 100.00
-ZECC w/o NL 98.44 8.33
-ZECC w/o VPS 92.55 100.00
-ZECC w/o ATE-VPS 99.49 1.45
-ZECC 18.91 100.00
-Table 3: The results of ablation study for components in NLBN.
+w/o HAE performs best by exploring environments greedily, but results in a higher TD.
+ZECC achieves a better balance by conducting both HAE and LAE simultaneously.
+To further illustrate this trade-off, we conducted an experiment by fixing TD for FBE in w/o ATE and ATE in ZECC.
+Once the agent reaches a TD threshold, it will conduct NLBN using partial estimated crowd distribution.
+The result is shown in Figure 5 (a).
+ZECC demonstrates better performance with less cost when TD is limited in most cases, illustrating that ZECC can efficiently estimates global crowd distribution.
+On the other hand, w/o ATE is not effective when TD is limited.
+NLBN.
+We then conducted ablation studies on NLBN and the results are presented in Table 5.
 w/o NLBN refers to using ATE results for crowd counting.
@@
 w/o ATE-VPS refers to not using the ATE-view-vector-based view point selection (ATE-VPS), but randomly select a navigation vector from the potential navigation vectors.
-Successful rate indicates the ratio of the reachable navigation point reported by the A* path planning algorithm.
-4.3 Ablation Study
-ATE.
-We conducted ablation studies by removing specific components from the ATE.
-The results are presented in Table 2.
-The configuration without ATE (w/o ATE) achieves the best performance because it explores the environments greedily and utilizes NLBN for precise detection, resulting in adequate perception.
-However, this approach leads to a larger temporal delay (TD).
-In contrast, ZECC achieves a better balance by conducting both HAE and LAE simultaneously.
-w/o HAE uses the same setting as w/o ATE.
-w/o LAE fails for it maintains at a high altitude and does not conduct close range detection, making it hard to capture targets.
-NLBN.
-We then conducted ablation studies on NLBN and the results are presented in Table 3.
-The findings indicate that omitting any component of NLBN results in a significant drop in either performance or success rates.
+Successful rate indicates the ratio of the reachable navigation point reported by path planning algorithm.
+The findings indicate that omitting any component of NLBN results in a significant drop in either performance or the success rate.
 Without NLBN, the absence of optimized viewpoints causes ZECC to revert to a ZSON method.
-When both NL and FBE-VPS are removed, most navigation points end up being located on obstacles, leading to a scarcity of reachable navigation points and incomplete exploration.
-Additionally, without VPS, the navigation points are positioned directly above the targets, causing the targets to fall out of the field of view, which results in corrupted observations.
+When both NL and FBE-VPS are removed, most navigation points end up being located on obstacles.
+Additionally, without VPS, the navigation points are positioned directly above the targets, causing the targets to fall out of the field of view.
+Table 4: The ablation study for ATE. ZECC achieves a better trade-off between performance and exploration efficiency (reducing 17% cost with 8% performance decline).
+Method MAPE (%) TD (m)
+w/o HAE 17.46 4633.67
+w/o LAE 88.08 1738.84
+ZECC 18.91 3804.63
+Table 5: The results of the ablation study for components in NLBN. ZECC achieves the best performance and success rate.
+Method MAPE (%) Successful rate (%)
+w/o NLBN 65.19 100.00
+w/o NL 98.44 8.33
+w/o VPS 92.55 100.00
+w/o ATE-VPS 99.49 1.45
+ZECC 18.91 100.00
 5.3 Hyper parameter Study
-We then examine the effects of parameters in the method, including GMM cluster size, navigation vector degree, navigation point range, and density map threshold.
+The influence of hyper parameters is studied, which include cluster size, navigation vector degree, navigation point range, and density map threshold.
 A gym-like scene featuring a densely packed crowd is utilized to test these parameters.
-The results are illustrated in Figure 6.
-GMM cluster size.
-As the size of the GMM cluster increases, performance tends to deteriorate.
-When the GMM cluster size is set too low, the method generates sparse navigation points, which prevents the agent from effectively exploring crowded areas, resulting in poor counting performance.
-Although a smaller GMM cluster size can initially yield high performance, the results become worse when the cluster sizes reach 30 and 40, respectively.
-Specifically, the cost for a cluster size of 30 is 4065.02 meters, while for a size of 40, it rises to 5125.44 meters.
-Opting for an extreme GMM cluster size may significantly increase costs without providing substantial improvements in performance.
-Therefore, it is crucial to carefully determine the GMM cluster size to find a balance between cost and performance.
+The results are illustrated in Figure 5 (b).
+Cluster size.
+Larger $\epsilon$ generates fewer observation with coarser detection, while low $\epsilon$ generates more navigation points and reduces efficiency (TD is 5125.44m for $\epsilon=30$, 4065.02m for $\epsilon=40$).
 Navigation vector degree.
-Performance is optimal at 20°, while it declines at other angles.
-This degradation occurs because the agent has a limited field of view; it cannot effectively observe targets outside this range.
-This highlights the importance of our navigation point generation method.
+The agent has a limited field of view for large $\zeta$ and is obstructed for low $\zeta$.
+This highlights the importance of the NLBN since the method generates robust navigation points for different scenes.
 Navigation point range.
-The performance is optimal at a distance of 8 meters, while performance declines at both shorter and longer ranges.
-A similar trend is observed with the navigation vector degree.
-At extremely close range, the agent’s field of view is restricted, and at long range, the agent is unable to gather detailed observations, which negatively impacts the subsequent target detection phase.
-Therefore, selecting the right viewpoint and generating appropriate navigation points are crucial for effective scene exploration algorithms.
+At close range, the agent’s field of view is restricted, and at long range, the agent is unable to gather detailed observations, which negatively impacts the subsequent target detection phase.
 Density map threshold.
-The results indicate that similar to GMM cluster size, performance declines as the threshold increases.
-A lower density map threshold results in a coarser estimation of the target zone, which in turn generates more navigation points.
-This allows for more detailed exploration and observation, leading to improved counting performance.
-Yet cost rises.
-For instance, when the density map threshold is 0.7, the travel distance is 4720.02 meters, and it increase to 7272.02 meters when the threshold is 0.4.
-This is a significant degradation in cost, yet improvement in performance is minor.
-This emphasizes the importance of parameter selection.
+A lower density map threshold leads to an expansion of the target area, resulting in an increased number of navigation points.
+However, this also raises the associated costs.
+For instance, when $\kappa=0.5$, TD is 6408.02 m, increasing to 7272.02 m when $\kappa=0.4$.
+This is a significant degradation in efficiency, yet the improvement in performance is not significant.
+Generally, the influence of hyper parameter is consist with tuition. ZECC provides effective performance if they are not set to extreme value.
+Figure 5: (a) Comparison of performance-cost trade-off.
+ZECC achieves a better trade-off when TD is limited.
+(b) The effect of four hyper parameters in ZECC.
+It shows that ZECC is effecive when the hyper-parameters are set in reasonable scopes.
+5.4 Case Study
+Figure 6: Navigation point (green) and trajectory (red) of different methods.
+Blcak dots are ground-truth.
+(a) ZECC.
+(b) OpenFMNAV.
+(c) MVC.
+ZECC shows less exploration in non-crowded areas while setting navigation points actively based on crowd distribution.
+Zoom in for a better view.
+We qualitatively illustrate how ZECC can alleviate occlusion and overlap while improving efficiency by comparing trajectory in two scenes with OpenFMNav and MVC-30.
+The results are shown in Figure 6.
+It shows that ZECC can effectively explore a complex occlusion environment, reduce exploration in low crowd density areas, and set navigation points based on crowd distribution.
+These features result in a trade-off between performance and cost.
 6 Conclusion
-In this study, Embodied Crowd Counting (ECC) is proposed, a task which enables interactive crowd counting.
-A simulator, the Embodied Crowd Counting Dataset (ECCD) is developed to enable related research for ECC.
-This dataset includes 60 diverse virtual environments with crowd density modeled by a prior probability distribution, approximating the reality.
+In this study, we propose a task that enables interactive crowd counting: Embodied Crowd Counting (ECC).
+A simulator, the Embodied Crowd Counting Dataset (ECCD), is developed to enable related research for ECC.
+This dataset includes 60 diverse virtual environments with crowd density modeled by a prior probability distribution, approximating reality.
 A method, Zero-shot Embodied Counting (ZECC), is proposed to verify this task.
-This is an active agent which can explore unknown environments without additional assistance.
+This is an active agent that can explore unknown environments without additional assistance.
 Active Top-down Exploration (ATE) is proposed to utilize Z-axis moving ability for exploration planning.
-This module is equipped with MLLM to active high altitude exploration (HAE) or low altitude exploration (LAE), balancing crowd counting performance and exploration cost.
-Normal-line based Navigation (NLBN) is proposed to select optimized navigation point for crowd observation.
-This module generates a navigation point from the top-down view and maintains an angle, alleviating overlap of the crowd.
-Simultaneously, the estimate navigation points enable obstacle avoidance and ensure that the crowd is in FOV.
+This module is equipped with MLLM to enable active high altitude exploration (HAE) or low altitude exploration (LAE), balancing crowd counting performance and exploration cost.
+Normal-line based Navigation (NLBN) is proposed to select an optimized navigation point for crowd observation.
+This module generates a navigation point from the top-down view and maintains an angle, alleviating the overlap of the crowd.
+Simultaneously, the estimated navigation points enable obstacle avoidance and ensure that the crowd is in FOV.
 Experiment results show that ZECC achieves a balance between performance and cost compared to recent navigation agents.
-Future research will focus on more efficient and scalable agents to enhance performance and adaptability.
+As the first work to propose ECC, we leave expending ECCD and ZECC to dynamic targets and real world application to future work, which are not considered by existing methods.
+7 Acknowledgment
+This work was supported by the Science and Technology Major Project of Jiangsu Province (No.BG2024041).
+This work was supported by the National Natural Science Foundation of China under Project 62406090.
+The work described in this paper was conducted in part by Dr ZHU Xinting, Jockey Club Global STEM Post-doctoral Fellow supported by The Hong Kong Jockey Club Charities Trust.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: We list the main contributions of this paper at the end of the introduction, and explain the scope and differences between our method and previous methods in both the abstract and the introduction.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We describe the limitations of our approach in the Conclusion section.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: We present comprehensive experimental results conducted on the dataset proposed, comparing our method with prior works to validate its effectiveness in Sec. 5.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: We describe the structure of our proposed model throughout in Sec.
+3 and implementation details of the experiment in Sec.
+4.
+To further ensure reproducibility, we will soon open source the code of our method.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: Once the paper is accepted, we will open source our code and the dataset.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: The implementation details are shown in Sec. 4.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: See experiment sections.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: The information is included in Sec. 4.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We have carefully read the NeurIPS Code of Ethics and have determined that our approach adheres to the relevant ethical guidelines.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: The broader impact is discussed in Sec. A.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: No such risk in this work.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: Airsim: https://paperswithcode.com/paper/airsim-high-fidelity-visual-and-physical
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: Our work does not release new assets. The data and models used in our work are publicly released.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: Our work does not involve human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: Our work does not involve human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [Yes]
+  79. Justification: We have carefully read the NeurIPS LLM policy and have determined that our approach adheres to the relevant ethical guidelines.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Appendix
+Appendix A Broader Impacts
+Further research and careful consideration are necessary when utilizing this technology, as the presented proposed method relies on the simulator, which may possess biases and could potentially result in negative societal impacts.
+Appendix B Visualization of ECCD
+Visualization of several scenes in ECCD is shown in Figure 7 and Figure 8.
+ECCD offers a diverse range of scenarios, encompassing both large-scale outdoor settings and indoor environments.
+It provides a hierarchical 3D structure, with crowds distributed across various positions within the 3D space, posing challenges for algorithms.
+Figure 7: Visualization of environments sampled from ECCD.
+Figure 8: 3D structure of samples from ECCD. Colorful dots represent crowd ground truth. Different colors represent different heights.
+Appendix C Details of ZECC
+C.1 ATE
+During HAE, the ATE receives RGB images captured from the bottom and surrounding areas of the agent.
+These images are organized into a panoramic view and transmitted to an MLLM in the form of a prompt.
+The image prompt and the text prompt are shown in Figure 9.
+The prompt utilizes the Chain of Thought mechanism to improve MLLM perception ability.
+Once the agent decides to switch to LAE, it will select a spare space directly under its horizontal location and conduct a landing operation.
+The pseudo-code is shown in Algorithm 1.
+Algorithm 1 Pseudo-Code of ATE
+1:Global distribution $D$, HAE map $M_{H}$, LAE map $M_{L}$, Prompt II
+2:
+[…]
+[…]
+[…]
+[…]
+[…]
+8:while not $done1$ do
+9:  $O_{t},p_{t}\leftarrow\text{getState}()$;
+10:  $s_{t}\leftarrow\text{MLLM}\left({{O_{t}};I}\right)$;
+11:  $M_{H}\leftarrow\text{updateMap}\left({{O_{t}};p_{t}}\right)$;
+12:  if $s_{t}>0.5$ then
+13:   $\text{toLAE}()$;
+14:   $done2\leftarrow False$;
+15:   while not $done2$ do
+16:     $O_{f},p_{f}\leftarrow\text{getState}()$;
+17:     $d_{f}\leftarrow\text{P}\left({\text{G}\left({{O_{f}}}\right),{p_{f}}}\right)$;
+18:     $M_{L}\leftarrow\text{updateMap}\left({{O_{f}};p_{f}}\right)$;
+19:     $D\leftarrow\text{updateDistribution}(D,d_{f})$;
+20:     $done2\leftarrow\text{toUnexplored}(M_{L})$;
+21:   end while
+22:   toHAE();
+23:  end if
+24:  $done1\leftarrow\text{toUnexplored}(M_{H})$;
+25:  $t\leftarrow t+1$;
+26:end while
+27:return $D$
+C.2 NLBN
+The method used to estimate the normal line of the crowd cluster plane is the Open3D package.
+The global crowd distribution is first divided into subareas using GMM cluster.
+Then, on each cluster, Open3D is used to estimate the normal line for each cluster.
+The normal lines are aligned with the vector (0,0,1).
+Then, the normal line at the cluster center is selected as the normal line of the cluster.
+Based on this representation, the navigation points can be obtained.
+The pseudo-code is shown in Algorithm 2.
+Algorithm 2 Pseudo-Code of NLBN
+1:Cluster size $\epsilon$, Navigation vector degree $\zeta$, Navigation point range $\eta$, Density map threshold $\kappa$, Global distribution $D$
+2:
+[…]
+[…]
+[…]
+6:for $i=1$ to NN do
+7:  ${\bf{d}}_{i}^{{\rm{cluster}}}\leftarrow\text{FitPlane}({\bf{x}}_{i}^{{\rm{cluster}}})$;
+8:  ${\bf{x}}_{i}^{{\rm{ATE}}}\leftarrow\text{getViewpoint}(D)$;
+9:  Compute $\{\mathbf{d}_{i1}^{\text{view}},...,\mathbf{d}_{im}^{\text{view}}\}_{i}$ using Eq. (4);
+10:  Compute ${\bf{d}}_{i}^{{\rm{ATE}}}$ using Eq. (5);
+11:  Compute ${\bf{d}}_{i}^{{\rm{view}}}$ using Eq. (6);
+12:  Compute ${\bf{x}}_{i}^{{\rm{view}}}$ using Eq. (7);
+13:end for
+14:while not donedone do
+15:  $done\leftarrow\text{toNextNaviPoint}(\{{\bf{x}}_{i}^{{\rm{view}}},...,{\bf{x}}_{N}^{{\rm{view}}}\})$
+16:end while
+C.3 FDC
+During FDC, the center of each detection box can be projected to the global crowd distribution by using the depth information.
+To re-identify targets, the space is divided into 3D voxels with size of 0.25 m.
+All the target with in a voxel is regared as the same target.
+This configuration is set for ZECC and all comparison methods.
+Figure 9: Prompt template used in ATE.
+Appendix D Occlusion analysis on ZECC
+In this section, we further test the ability of ZECC to alleviate occlusion in the complex environments.
+For each ground truth person, we first select the navigation point which has the minimal Euclidean distance to it, and then test whether occlusion occurs between the navigation point and the person.
+This is implemented by first projecting the global point cloud to the vector from navigation point to the person.
+Then, if the minimal distance between the global point cloud and the projected points is lower than a threshold, the person is obstructed.
+To formulate this, the unit vector from a person to its nearest navigation point is calculated as:
+${\bf{v}}_{i}^{{\rm{navi}}}=\frac{{{\bf{P}}_{i}^{{\rm{navi}}}-{\bf{P}}_{i}^{{\rm{crowd}}}}}{{\left\|{{\bf{P}}_{i}^{{\rm{navi}}}-{\bf{P}}_{i}^{{\rm{crowd}}}}\right\|}},$ (8)
+where ${\bf{P}}_{i}^{{\rm{navi}}}$} is coordinate of the navigation point and ${\bf{P}}_{i}^{{\rm{crowd}}}$} is the coordinate of the person. The vector from a global point cloud to the person is:
+${\bf{v}}_{ij}^{{\rm{global}}}={\bf{P}}_{i}^{{\rm{navi}}}-{\bf{P}}_{j}^{{\rm{global}}},$ (9)
+where ${\bf{P}}_{j}^{{\rm{global}}}$ is a random coordinate from the global point cloud. The projected length is:
+${l_{ij}}={\bf{v}}_{ij}^{{\rm{global}}}\cdot{\bf{v}}_{i}^{{\rm{navi}}},$ (10)
+and the projected point is:
+${\bf{P}}_{j}^{{\rm{proj}}}={\bf{P}}_{i}^{{\rm{crowd}}}+{l_{ij}}*{\bf{v}}_{i}^{{\rm{navi}}}\text{ s.t. }{l_{ij}}\geq 0\wedge{l_{ij}}\leq\left\|{{\bf{v}}_{ij}^{{\rm{global}}}}\right\|.$ (11)
+The condition of determining the person is obstructed is:
+$\mathop{{\rm{min}}}\limits_{j}\left\|{{\bf{P}}_{j}^{proj}-{\bf{P}}_{j}^{global}}\right\|\leq\lambda.$ (12)
+In the experiment, $\lambda$ is set to 0.5 m.
+We compare ZECC with OpenFMNav and MVC-30.
+The ratio between the number of obstructed person and the number of ground truth person of the three methods are shown in Table 7.
+It shows that ZECC suffers from the less occlusion, which benefits the crowd counting results.
+Appendix E Failure case study
+ZECC fails when MLLM makes wrong decision.
+This is mainly due to occlusion by buildings.
+Figure 10 shows an example of MLLM planning failure, where the agent does not choose to conduct LAE at a crowd area.
+Figure 10: A case of ZECC failure. (a) Agent trajectory. (b) Top-down RGB. (c) MLLM reasoning. Blue box: ground truth crowd. Because of occlusion, MLLM failed to perceive the crowd during HAE.
+Appendix F Quantitative results in real environment
+To test ZECC in the real environment, we use two outdoor crowd scene to verify the proposed method.
+The drone model used is the Taobotics Q300.
+The captured images are transmitted back to a ground server using a remote communication protocol and fed into VGGT [49] to estimate relative point clouds and poses.
+The proposed NLBN is then employed to calculate relative navigation points.
+The drone’s absolute geographic coordinates and poses are recorded in real-time by its GPS.
+By leveraging the relationships between absolute and relative geographic coordinates and poses, the navigation points calculated by NLBN are mapped to absolute navigation points.
+After the ground server controls the drone to fly to these absolute navigation points, it captures images of the crowd.
+We utilize Grounding DINO and the crowd counting model Generalized Loss to perform crowd detection or counting on images captured from both distant and close-up perspectives, respectively.
+The ground truth for crowd annotations is manually labeled.
+The visualization of the drone observation before and after adjusting its position to the NLBN navigation point is shown in Figure 11.
+The quantitative results are shown in Table 7.
+Before applying NLBN, the drone can not obtain a clear view of the crowds in the target area.
+The GL estimation result is fuzzy, and due to crowd overlap, people in the back rows are obstructed by those in the front rows.
+This results in the corrupted detection of GD.
+By adjusting to the NLBN navigation point, the drone is able to observe the target area from a top-down perspective, alleviating occlusion of people in back rows and improving counting performance significantly.
+This demonstrates how NLBN addresses occlusion.
+Table 6: Obstructed person ratio.
+Method Ratio (%)
+OpenFMNav 42.81
+MVC-30 21.77
+ZECC 4.52
+Table 7: Performance of real scenarios.
+Methods MAPE (%)
+Scenario 1 Scenario 2
+w/o NLBN + GL 41.41 57.76
+w/o NLBN + GD 73.74 78.44
+NLBN + GL 22.50 29.31
+NLBN + GD 15.17 19.83
+Figure 11: Quantitative results in real environment.
+(a), (d) Output of GL.
+The marked areas indicate the target zone.
+(b), (e) Point cloud of the scene.
+Red lines: candidate view vector.
+Green lines: the drone view vector.
+Orange lines: selected view vector.
+(c), (f) GD detection results.
+Red boxes: detected crowds.
+Green boxes: ground truth crowds.
+First row: before applying NLBN.
+Second row: after applying NLBN.
+(a), (b) and (c) are Scenario 1 and (d), (e) and (f) are Scenario 2.
+By adjusting the camera position to the NLBN navigation point, the target detection result is improved.